OpenStudy (anonymous):

Could someone please explain to me how to get the answer in fraction form? Use the quadratic formula to solve the equation. –4x2 – 3x + 2 = 0

OpenStudy (anonymous):

wont factor using integers

OpenStudy (anonymous):

make your life easier by starting with \[4x^2-3x-2=0\] then compute \[\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] with \[a=4,b=3,c=-2\]